I'm sorry, but I miswrote the problem. Thanks again!
The question is:
Solve for y:
y+10/y+3 = y+1/y-1

y = (?)

Answer by nadinadan(32)

Let's try again:
y + 10/y +3 = y +1/y -1
Same denominator so we multiply the first number by y, the third by y
The fourth by y and the 6th by y.
So we have y^2+10+3y-y^2-1+y=0 Do the operations
4y+9=0 so y=-9/4
Hope we got it this time.